# NocPlace: Nocturnal Visual Place Recognition Using Generative and Inherited Knowledge Transfer

### 27 Feb 2024 v1

@@ 4 Methodology
Figure 4: Schematic Diagrams of NocPlace. In the training phase, we transfer the night knowledge into the training data and use the existing label information. We inherit the descriptors extracted from the corresponding data by the pre-trained daytime VPR model and use them to supervise the NocPlace model. During the test phase, NocPlace extract features from the query image and query them in the descriptor database extracted by the daytime VPR model.
In this section, we detail the training and testing strategies of NocPlace, which encompass generative knowledge transfer, inherited knowledge transfer, and a divide-and-conquer retrieval approach. The upper part of Fig. 4 corresponds to Sec. 4.1 and 4.2, and the lower part corresponds to Sec. 4.3.
@@ 4.1 Generative Knowledge Transfer
Employing the VPR training framework [9, 11] with Large Margin Cosine (LMC) loss [55], we can either fine-tune the model using both source and generated domain data or train exclusively with the generated domain data. Theoretically, the former approach balances performance between daytime and nighttime, while the latter emphasizes nighttime performance.
@@ 4.2 Inherited Knowledge Transfer
The advantage of LMC loss over triplet loss is that it avoids mining difficult samples, which makes it easier to achieve the current state-of-the-art results on large-scale training data sets [9, 11]. However, LMC loss also has the disadvantage that it cannot explicitly construct cross-domain contrastive relationships. Meanwhile, we assume that a pre-trained VPR model has learned a representation suitable for daytime VPR. Therefore, we hope to minimize the distribution difference between the day and night domains to inherit the representation ability of the pre-trained model. Specifically, we save the global descriptors extracted from daytime images by the pre-trained model, align them with the nighttime images and minimize the network outputs by Kullback–Leibler (KL) divergence.
@@ 4.3 Retrieval based on Partial Divide-and-Conquer
We have decomposed VPR into two distinct sub-problems, Day-to-Day VPR and Night-to-Day VPR, addressing these challenges using two separate models. This approach is novel, and we explore its practicality and potential for optimization.
In real-world applications, computers can easily differentiate between day and night using straightforward methods that require only one-time operations. For instance, we can determine day and night simply by the system time of a device. Therefore, NocPlace involves only model parameter switching within the same architecture, without introducing additional inference overhead. This significantly enhances performance in nighttime environments.
An inherent limitation of NocPlace is the requirement for computers to store two models and two databases. With the advent of an edge-cloud collaborative VPR system and the ease of scalable storage solutions in modern computing, this drawback becomes a manageable trade-off. However, if both the database and the model are on the edge, such as in decentralized autonomous robots, it would significantly consume storage resources. For instance, each image in the SF-XL test set is output as a feature vector of 512 dimensions, and the entire database size would be a staggering 5.7 GB. To overcome this challenge, we only use NocPlace to process the query image during the testing phase, while the database retains the features extracted by the pre-trained model. This design also aligns with the design of IKT.
@@ 5 Experiments
In this section, we begin by providing a description of the implementation details, the test sets, and the evaluation metrics. Subsequently, we present a comprehensive range of quantitative experiments focusing on night-to-day visual place recognition, night-to-day visual localization, and day-to-night image-to-image translation.
Dataset Name SVOX SF-XL MSLS Tokyo 24/7 Tokyo 24/7 Aachen v1.1 RobotCar Seasons v2 Night Night Night Night Sunset Night Night # queries 823 466 55 105 105 192 438 # database 17k 2.8M 18.9k 76k 76k 6.7k 6.9k Orientation frontal-view multi-view frontal-view multi-view multi-view multi-view frontal-view Scenery mostly urban urban urban urban urban mostly urban mostly urban Acquisition Method vehicles handheld devices vehicles handheld devices handheld devices handheld devices vehicles
Table 1: Overview of night VPR and VL test sets. We can see that the test sets listed above includes various scales, scenarios, collection methods, and database image orientations.
@@ 5.1 Implementation details
5.1.1 Night-to-Day Visual Place Recognition
Architecture. The core concept of NocPlace revolves around transferring generated nocturnal information and existing day information into a new model. Therefore, it is expected to be compatible with all previous methods and models, enhancing their performance specifically in nocturnal scenarios. In order to illustrate the potential and scalability of NocPlace, we followed a methodology akin to CosPlace [9] and EigenPlaces [11], leveraging a simple convolutional neural network architecture, specifically VGG-16 [48] and ResNet-50 [25]. The architecture was connected to a GeM pooling layer [41] and a fully connected layer, with the dimensionality of the descriptors produced by NocPlace being determined by the specifications of the fully connected layer.
Training. NocPlace employed CosPlace (512 dimensions) or EigenPlaces (2048 dimensions) as pre-trained models, requiring only 1 epoch of fine-tuning on SF-XL-NC with a learning rate of 1e-6. An important distinction is that NocPlace no longer utilizes any data augmentation techniques. The parameter $\alpha$ in the loss function can be adjusted by observing the initial loss of $L_{\mathrm{LMC}}$, but it does not require meticulous tuning. For instance, in training a model based on the ResNet-50 with an output dimension of 512, this parameter was set to 30.
5.1.2 Night-to-Day Visual Localization
@@
5.1.3 Day-to-Night Image-to-Image Translation
We did not modify any training parameters and followed the official tutorial directly [38, 57]. It is worth noting the difference in image resolution between NightCities and the VPR training sets. To address this, we proportionally resize the NightCities images, keeping the shorter side at 640 pixels, and subsequently apply random cropping to achieve a size of 512$\times$512 during the training process. This approach ensures as much resolution consistency as possible between the I2I and VPR training sets. During the evaluation, we trained NEGCUT on multiple data sets mentioned in Sec. 5 and used NightStreets as the test set.
@@ 5.2 Test sets
We conducted experiments on multiple (6) nighttime VPR datasets, which encompass challenges frequently discussed in daytime VPR scenarios: large-scale, viewpoint variations, scale changes, etc. Some specific details are presented in Tab. 1. Aside from adaptive modifications to the Tokyo 24/7 [5] and MSLS [58] datasets, the rest follow prior works [45, 65, 46, 11].
Tokyo 24/7 [5] is highly challenging in terms of scale, with a vast database size and a three-year gap between query images and database images. Notably, the query images for each location are distributed across three time periods: daytime, sunset, and nighttime. To the best of our knowledge, few studies [6, 5] segment this test set according to these time frames, somewhat overshadowing the impact of nighttime.
MSLS [58] is a crowdsourced dataset derived from dashcam recordings. We followed the validation set configuration from [9] and filtered out nighttime images. Although our experiment approach differs from [11] and other studies, it ensures a fair comparison in our experiments.
@@
NetVLAD [5] (CVPR’16)
SFRS [19] (ECCV’20)
CAE-VPR [63] (Robotica’23)
CosPlace [9] (CVPR’22)
EigenPlaces [11](ICCV’23)
@@
CosPlace [9](CVPR’22)
Conv-AP [1] (NC’22)
MixVPR [2] (WACV’23)
EigenPlaces [11] (ICCV’23)
@@
CosPlace [9](CVPR’22)
Conv-AP [1] (NC’22)
EigenPlaces [11] (ICCV’23)
[11] + ToDay-GAN [4] (ICRA’19)
@@
Table 2: Recall@1 / 5 / 10 on Night-to-Day VPR datasets. The best results under the same model and the same data set are bolded.
@@
Table 3: Localization success rate at different localization thresholds, recall@N, and datasets. [50]. The localization thresholds are (0.25m, 2°) / (0.5m, 5°) / (5m, 10°). Specifically, recall numbers are set to 1, 5, 10. The best results for the same data set and the same threshold are bolded.
5.3 Evaluation metrics
Night-to-Day Visual Localization. We evaluate VPR through the localization success rate under varying recall@N metrics. It is crucial to note that the visual localization task demands more rigorous precision than conventional VPR benchmarks.
Day-to-Night Image-to-Image Translation. Our evaluation of translated images encompasses two primary facets: realism and faithfulness [71, 38, 57]. Realism quantitatively assesses the disparity between the target and translated image distributions. Faithfulness gauges the extent of congruence between an original image and its translation on an individual basis. For a precise measure of realism, we leverage the established Frechet Inception Score (FID). In measuring Faithfulness, we utilize metrics like the pixel-wise $L_{2}$ distance, peak-signal-to-noise ratio (PSNR), and the structural similarity index measure (SSIM) for each paired input-output.
5.4 Comparison with previous work
We have conducted extensive comparisons of numerous methods on various datasets, an endeavor unprecedented in nighttime VPR, as shown in Tab. 2 and Tab. 3. These methods range from the classical NetVLAD[5], To-Day GAN [4] to more recent contributions over the past three years, including SFRS [19], CosPlace [9], MixVPR [2], CAE-VPR [63], Conv-AP [3], and EigenPlaces [11]. We replicated their open-source models and code for our experiments. It is worth noting that these methods were also tested on the SVOX Night dataset in [11] and our results are consistent with theirs. The experimental results can be distilled into the following key observations:
  • NocPlace is the undisputed winner in nighttime scenarios. Its performance enhancements are substantial, even when compared to EigenPlaces (ICCV’2023).
  • The sunset scenes in Tokyo 24/7 present a mix of numerous daytime and fewer nighttime images, which exposes the degradation of NocPlace during daylight conditions.
  • On both SVOX Night and Tokyo 24/7 Night datasets, NocPlace achieves a significant improvement in cross-architectural performance.
  • The advantages of NocPlace are magnified on the MSLS Night dataset, due to the limited number of query images (less than 100) and redundancy (sequential frames).
  • In scenarios with particularly large databases, such as the SF-XL dataset, the improvements of NocPlace are objectively manifested. Conversely, in smaller databases like the Aachen v1.1 dataset where the VPR challenge is reduced, the gains of NocPlace become less pronounced, even trailing behind methods with higher feature dimensions, such as SFRS.
  • On the RobotCar Seasons v2 Night dataset, CAE-VPR achieved the best results, and ToDay GAN was able to enhance the performance of EigenPlaces. However, they showed poor generalization on other datasets, which confirms the viewpoint we stated earlier. Additionally, it is noteworthy that the training set chosen by CAE-VPR and the test set for VL seem to have an intersection (Their division methods are inconsistent).
@@ 5.5 Ablations
Training set IKT OD Tokyo 24/7 SVOX Day Night Night Night 80.0 / 88.6 / 92.4 51.6 / 68.8 / 76.1 83.8 / 94.3 / 94.3 62.3 / 77.4 / 81.9 83.8 / 93.3 / 96.2 63.2 / 78.7 / 84.8 89.5 / 93.3 / 94.3 63.9 / 77.9 / 84.9 86.7 / 92.4 / 92.4 65.6 / 80.8 / 85.3 87.6 / 93.3 / 92.4 66.6 / 80.9 / 87.1 91.4 / 93.3 / 93.3 67.3 / 81.0 / 86.6
Table 4: Ablation on designs, shown as the Recall@ 1/5/10 obtained with a ResNet-50 with output dimensionality 512 on two datasets.
We conducted ablation experiments on the ResNet-50 architecture with an output dimension of 512 for NocPlace, as shown in Tab. 4. It is evident that generating nighttime data and incorporating the IKT loss both result in significant performance improvements. Here, OD means using the original database, which also means the partial divide-and-conquer based retrieval mentioned before. By using OD, we sacrificed accuracy to maintain the practicality of the method. This implies that without using OD, NocPlace in Tab. 2 and Tab. 3 could achieve even better experimental results.
@@ 5.6 Comparision with previous datasets
Dataset FID $\downarrow$ L2 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ Night2Day 243.90 164.79 8.74 0.21 LOL 99.62 191.99 7.40 0.14 RobotCar 148.88 156.29 9.29 0.32 NightStreets 63.54 121.62 11.75 0.60 NightCities 62.21 98.12 13.98 0.75
Table 5: Quantitative comparison of generated results for different Night-to-Day datasets. The best results are shown in bold.
We focus on the quality of the day-to-night dataset. To provide an objective assessment regarding the dataset in our work, we trained NEG-CUT on the datasets previously mentioned and on the dataset we proposed. Only the Night2Day dataset uses the pre-trained model of Pix2Pix [27] because it is a paired day-night dataset. The test set directly utilized NightStreets. As shown in the Tab. 5, the NightCities model achieved the best results on evaluation metrics commonly used in the I2I translation domain, even surpassing the NightStreets model.

### 21 Mar 2024 v2

@@ 4 Methodology
In this section, we detail the training and testing strategies of NocPlace, which encompass generative knowledge transfer, inherited knowledge transfer, and a divide-and-conquer retrieval approach. The upper part of Fig. 4 corresponds to Sec. 4.1 and Sec. 4.2, and the lower part corresponds to Sec. 4.3.
Figure 4: Schematic Diagrams of NocPlace. In the training phase, we transfer the night knowledge into the training data and use the existing label information. We inherit the descriptors extracted from the corresponding data by the pre-trained daytime VPR model and use them to supervise the NocPlace model. During the test phase, NocPlace extracts features from the query images and queries them in the descriptor database extracted by the daytime VPR model.
@@ 4.1 Generative Knowledge Transfer
Employing the VPR training framework [8, 11] with Large Margin Cosine (LMC) loss [71], we can either fine-tune the model using both source and generated domain data or train exclusively with the generated domain data. The former approach balances performance between daytime and nighttime, while the latter emphasizes nighttime performance.
@@ 4.2 Inherited Knowledge Transfer
The advantage of LMC loss over triplet loss is that it avoids mining difficult samples, which makes it easier to achieve the current state-of-the-art results on large-scale training data sets [8, 11]. However, LMC loss also has the disadvantage that it cannot explicitly construct cross-domain contrastive relationships. Meanwhile, we assume a pre-trained VPR model has learned a representation suitable for daytime VPR. Therefore, we aim to minimize the distribution difference between the day and night domains to inherit the representation ability of the pre-trained model. Specifically, we save the global descriptors extracted from daytime images by the pre-trained model, align them with the nighttime images, and minimize the network outputs by Kullback–Leibler (KL) divergence.
@@ 4.3 Retrieval based on Partial Divide-and-Conquer
We have decomposed general VPR into two distinct sub-problems, Day-to-Day VPR and Night-to-Day VPR, addressing these challenges using two separate models. This approach is novel, and we explore its practicality and potential for optimization.
In real-world applications, computers can easily differentiate between day and night using straightforward methods that require only one-time operations. For instance, we can determine day and night simply by the system time of a device and approximate latitude and longitude information  [^2] https://www.timeanddate.com/sun/italy/milan. Therefore, NocPlace involves only model parameter switching within the same architecture, without introducing additional inference overhead.
An inherent limitation of NocPlace is the requirement for computers to store two models and two databases. With the advent of an edge-cloud collaborative VPR system and the ease of scalable storage solutions in modern computing, this drawback becomes a manageable trade-off. However, if both the database and the model are on the edge, such as in decentralized autonomous robots, it would significantly consume storage resources. For instance, each image in the SF-XL test set is output as a feature vector of 512 dimensions, and the entire database size would be a staggering 5.7 GB. To overcome this challenge, we only use NocPlace to process the query image during the testing phase, while the database retains the features extracted by the pre-trained model. This section corresponds to the experimental setup in Tab. 4 regarding whether the original database is used.
@@ 5 Experiments
In this section, we describe the implementation details, the test sets, and the evaluation metrics. Subsequently, we present a comprehensive range of quantitative experiments focusing on night-to-day visual place recognition, night-to-day visual localization (VL), and day-to-night image-to-image translation. Finally, extensive qualitative experimental results are presented in the supplementary material.
@@ 5.1 Implementation details
The architecture of NocPlace. The core idea of NocPlace revolves around transferring generated nocturnal knowledge and inherited day knowledge into existing models. Therefore, it is expected to be compatible with all previous methods and models, enhancing their performance in nocturnal scenarios. To illustrate the potential and scalability of NocPlace, we followed a methodology akin to CosPlace [8] and EigenPlaces [11], leveraging a simple convolutional neural network architecture, specifically VGG-16 [64] and ResNet-50 [29]. The architecture was connected to a GeM pooling layer [54] and a fully connected layer, with the dimensionality of the descriptors produced by NocPlace being determined by the specifications of the fully connected layer.
The Training of NocPlace. NocPlace employed CosPlace (512 dimensions) or EigenPlaces (2048 dimensions) as pre-trained models, requiring only one epoch of fine-tuning on SF-XL-NC with a learning rate of 1e-6. An important distinction is that NocPlace no longer utilizes any data augmentation techniques. The parameter $\alpha$ in the loss function can be adjusted by observing the initial loss of $L_{\mathrm{LMC}}$, but it does not require meticulous tuning. For instance, in training a model based on the ResNet-50 with an output dimension of 512, this parameter was set to 30.
5.1.1 Night-to-Day Visual Localization.
@@
5.1.2 Day-to-Night Image-to-Image Translation.
We did not modify any training parameters and followed the official tutorial directly [50, 73]. The difference in image resolution between NightCities and the VPR training sets is worth noting. To address this, we proportionally resize the NightCities images, keeping the shorter side at 640 pixels, and subsequently apply random cropping to achieve a size of 512$\times$512 during the training process. This approach ensures as much resolution consistency as possible between the I2I and VPR training sets. During the evaluation, we trained NEGCUT on multiple data sets mentioned in Sec. 5 and used NightStreets as the test set.
@@ 5.2 Test sets
Table 1: Overview of night VPR and VL test sets. We can see that the test sets listed below includes various scales, scenarios, collection methods, and orientations.
Dataset Name SVOX SF-XL MSLS Tokyo 24/7 Tokyo 24/7 Aachen v1.1 RobotCar Seasons v2
Night [12] Night [6] Night [74] Night [5] Sunset [5] Night [55] Night [43]
# queries 823 466 55 105 105 192 438
# database 17k 2.8M 18.9k 76k 76k 6.7k 6.9k
Orientation frontal-view multi-view frontal-view multi-view multi-view multi-view frontal-view
Scenery mostly urban urban urban urban urban mostly urban mostly urban
Acquisition Method vehicles handheld devices vehicles handheld devices handheld devices handheld devices vehicles
We conducted experiments on multiple nighttime VPR datasets, encompassing challenges frequently discussed in daytime VPR scenarios: large-scale, viewpoint variations, and scale changes. Some details are presented in Tab. 1. Aside from reasonable modifications to Tokyo 24/7 [5] and MSLS [74], the rest follow prior works [59, 83, 58, 11].
Tokyo 24/7 [5] is highly challenging in scale and has a three-year gap between query images and database images. Notably, the query images for each location are distributed across three time periods: daytime, sunset, and nighttime. To the best of our knowledge, few studies [6, 5] segment this test set according to these time frames, somewhat overshadowing the impact of nighttime.
MSLS [74] is a crowdsourced dataset derived from dashcam recordings. We followed the validation set configuration from [8] and filtered out nighttime images. The experimental results for MSLS Night and the results on the public MSLS leaderboard [74] show significant differences, but the comparison is fair.
5.3 Evaluation metrics
Table 2: Recall@1 / 5 / 10 on Night-to-Day VPR datasets. The best results under the same model and the same data set are shown in red.
@@ 5.3 Evaluation metrics
NetVLAD [5]
SFRS [22]
CAE-VPR [80]
CosPlace [8]
EigenPlaces [11]
@@
CosPlace [8]
Conv-AP [1]
MixVPR [2]
EigenPlaces [11]
@@
CosPlace [8]
Conv-AP [1]
EigenPlaces [11]
[11] + ToDay GAN [4]
@@
Night-to-Day Visual Localization. We evaluate VPR through the localization success rate under different recall@N metrics [55, 38, 58, 66]. It is crucial to note that the VL metrics demand more rigorous precision than VPR metrics.
Day-to-Night Image-to-Image Translation. Our evaluation of translated images encompasses two primary facets: realism and faithfulness [89, 50, 73]. Realism quantitatively assesses the disparity between the target and translated image distributions. Faithfulness gauges the extent of congruence between an original image and its translation. We leverage the established Frechet Inception Score (FID) for a precise measure of realism. In measuring Faithfulness, we utilize metrics like the pixel-wise $L_{2}$ distance, peak-signal-to-noise ratio (PSNR), and the structural similarity index measure (SSIM) for each paired input-output.
5.4 Comparison with previous work
We have conducted extensive comparisons of numerous methods on various datasets, an endeavor unprecedented in nighttime VPR, as shown in Tab. 2 and Tab. 3. These methods range from the classical NetVLAD[5], To-Day GAN [4] to more recent contributions over the past three years, including SFRS [22], CosPlace [8], MixVPR [2], CAE-VPR [80], Conv-AP [3], and EigenPlaces [11]. We replicated their open-source models and code for our experiments. It is worth noting that these methods were also tested on SVOX Night in [11], and our results are consistent with theirs. The experimental results can be distilled into the following key observations:
  • NocPlace is the undisputed winner in nighttime scenarios. Its performance enhancements are substantial, even compared to EigenPlaces (ICCV’2023).
  • SF-XL Night (CVPRW’2023) is the most challenging nighttime VPR test set. On this test set, the Recall@1 of NocPlace is 4.7% higher than the Recall@1 of EigenPlaces.
  • On both SVOX Night and Tokyo 24/7 Night, NocPlace even shows performance improvements across architectures and feature dimensions. For example, VGG16-based NocPlace excels over ResNet50-based EigenPlaces.
  • The advantages of NocPlace are magnified on MSLS Night due to the limited number of query images (less than 100) and redundancy (sequential frames).
  • The sunset scenes in Tokyo 24/7 present a mix of numerous daytime and fewer nighttime images, which exposes the degradation of NocPlace during daylight conditions. As long as model switching can be considered as described in Sec. 4.3, this degradation will not occur in practice.
  • In the case of high-precision positioning and small-scale databases, the challenge and importance of VPR are weakened. On Aachen v1.1, the advantage of NocPlace became less clear, trailing slightly behind SFRS (ECCV’2020) by 1%.
  • On RobotCar Seasons v2 (RS-v2), ToDay GAN enhanced the performance of EigenPlace, and CAE-VPR achieved the best performance. However, they show poor generalization on other test sets, which supports the viewpoints made in Sec. 1 and Sec. 2. In particular, ToDay GAN needs to execute two models in series when performing a VPR task, which consumes additional computing resources. RS-v2 and the training set of CAE-VPR are the same part of the original RobotCar. So, the results of CAE-VPR may be unfair to NocPlace and other methods on RS-v2.
Table 3: Localization success rate at different localization thresholds, recall@N, and datasets. [66]. The localization thresholds are (0.25m, 2°) / (0.5m, 5°) / (5m, 10°). Specifically, recall numbers are set to 1, 5, 10. The best results for the same data set and the same threshold are shown in red.
@@ 5.5 Ablations
Table 4: Ablations. This table shows the Recall@1/5/10 obtained with a ResNet-50 with output dimensionality 512 on two datasets. IKT means Inherited Knowledge Transfer, and OD means using the original database.
Training set IKT OD Tokyo 24/7 SVOX
Day Night Night Night
80.0 / 88.6 / 92.4 51.6 / 68.8 / 76.1
83.8 / 94.3 / 94.3 62.3 / 77.4 / 81.9
83.8 / 93.3 / 96.2 63.2 / 78.7 / 84.8
89.5 / 93.3 / 94.3 63.9 / 77.9 / 84.9
86.7 / 92.4 / 92.4 65.6 / 80.8 / 85.3
87.6 / 93.3 / 92.4 66.6 / 80.9 / 87.1
91.4 / 93.3 / 93.3 67.3 / 81.0 / 86.6
We conducted ablation experiments on the ResNet-50 architecture with an output dimension of 512 for NocPlace, as shown in Tab. 4. It is evident that generating nighttime data and incorporating the IKT loss results in significant performance improvements. Here, OD means using the original database, which also means the partial divide-and-conquer-based retrieval mentioned before. By using OD, we sacrificed accuracy to maintain the practicality of NocPlace. This implies that without using OD, NocPlace in Tab. 2 and Tab. 3 could achieve even better experimental results.
@@ 5.6 Comparision with previous datasets
Table 5: Quantitative comparison of NEG-CUT [73] on different night-to-day datasets. The best results are shown in red. Only the day2night dataset uses the pre-trained model of Pix2Pix [31].
Dataset Method FID $\downarrow$ L2 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
day2night [36] Pix2Pix 243.90 164.79 8.74 0.21
LOL [75] NEG-CUT 99.62 191.99 7.40 0.14
RobotCar [43] NEG-CUT 148.88 156.29 9.29 0.32
NightStreets NEG-CUT 63.54 121.62 11.75 0.60
NightCities (Ours) NEG-CUT 62.21 98.12 13.98 0.75
We focus on the quality of day-to-night datasets and the performance of image-to-image methods. We discuss the former here and the latter in the supplementary material. To objectively assess these datasets, we trained NEG-CUT on the previously mentioned and proposed datasets. Only the day2night dataset [36] uses the pre-trained model of Pix2Pix [31] because it is a paired day-night dataset. The test set uses NightStreets to comply with the principle of single-variable experiments. As shown in the Tab. 5, the model trained on NightCities achieved the best evaluation metrics commonly used in the I2I translation domain, even surpassing the model trained on NightStreets. Considering that NightStreets is reorganized from the VPR test set, these results also show that the model trained on NightCities can generate data that is more similar to the test data.